OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (mathmate):

Are you familiar with the equation of the line through two points? (y-y1)= m(x-x1) where m is the slope and m=(y2-y1)/(x2-x1) The choices are given in the form of the above equation. So by applying the equation, you would be able to make your choice.

OpenStudy (mathmate):

Hint: (−3, 2) is your point (x1,y1)

OpenStudy (anonymous):

i did it that way and for some reason its not coming out to be none of the answers above so do you think you can exolain it some more

OpenStudy (mathmate):

(−3, 2) is (x1,y1), so x1=-3, y1=2. Substitute in the above equation: y-2 = m(x-(-3)) You should be able to make your choice already! A better way is to confirm the value of m.

OpenStudy (anonymous):

is it c

OpenStudy (mathmate):

y-2 = m(x-(-3)) reduces to y-2 = m (x+3) Is that clearer?

OpenStudy (anonymous):

yes thank you very much

OpenStudy (mathmate):

Great, good job!